Determine whether each set contains measures of the side lengths of a triangle?
A.(2,3,1) B. (2,3.4) C. (2,2,3) D. (18,30,24)
and which of sets contains side lengths of a triangle?

Answer by CharlesG2(834)

"Determine whether each set contains measures of the side lengths of a triangle?
A.(2,3,1) B. (2,3.4) C. (2,2,3) D. (18,30,24)
and which of sets contains side lengths of a triangle?"

the sum of any 2 sides of a triangle must be greater than the measure of the 3rd side
A.(2,3,1):
2 + 3 = 5, 5 > 1? yes
3 + 1 = 4, 4 > 2? yes
2 + 1 = 3, 3 > 3? no
not a triangle
B. (2,3.4):
assuming this is (2,3,4) and not (2,3.4)
2 + 3 = 5, 5 > 4? yes
3 + 4 = 7, 7 > 2? yes
2 + 4 = 6, 6 > 3? yes
this is a triangle
if this was (2,3.4) then we only know 2 sides and do not have enough information to determine if this is a triangle
C. (2,2,3):
2 + 2 = 4, 4 > 3? yes
2 + 3 = 5, 5 > 2? yes
2 + 3 = 5, 5 > 2? yes
this is a triangle
D. (18,30,24):
18 + 30 = 48, 48 > 24? yes
30 + 24 = 54, 54 > 18? yes
18 + 24 = 42, 42 > 30? yes
this is a triangle